An operations manager, after monitoring the production forecast for the last 12 months, calculated the following errors: MAD=34.7; MSE=127.6; MAPE=12.5%. She then collected actual vs forecast data for an additional 8 months:

Period 1 2 3 4 5 6 7 8
Actual - Forecast 30 27 -25 13 -31 5 -12 18

A) Calculate 3 sigma upper and lower control limits for the data

B)using these limits is most recent forecast "in-control"?

C)should she consider changing the forecasting methodology?
Answer #1

Solution:-

Given data:-

MAD=34.7

MSE=127.6

MAPE=12.5%

(a)Calculate 3 sigma upper and lower control limits for the data?

3 sigma upper control limit = 0+3 sqrt(MSE)

Where MSE=127.6

3 sigma upper control limit = 0+3 sqrt(127.6)

3 sigma upper control limit = 33.88805099

3 sigma lower control limit = 0-3sqrt(MSE)

Where MSE=127.6

3 sigma lower control limit = 0-3sqrt(127.6)

3 sigma lower control limit = -33.88805099

(b)Using these limits is most recent forecast "in-control"?

We can see forecast errors for additional 8 months are within control limits.

(c)Should she consider changing the forecasting methodology?

No she should not consider changing the forecasting methodology